OpenStudy (anonymous):

Write the equation of the line that is perpendicular to the line y =3 over 2x+4 and passes through the point (-6,3) A. y=3 over 2x-1 B. y=3 over 2x+3 C. y=-2 over 3x-1 D. y=-2 over 3x+3

OpenStudy (anonymous):

@iluvreading

OpenStudy (anonymous):

What do you mean by over 2x+4?

OpenStudy (anonymous):

as in a fraction

OpenStudy (anonymous):

you mean y=3/2x+4?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

I think there is some problem with your question. I know about y=2x+4/3, but i havent done anything like where a xy term comes. I'm afraid i wouldnt be able to help you.

OpenStudy (anonymous):

poo

OpenStudy (anonymous):

Oh dang. I get it. Your question is this. y=(3/2)x+4 right?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Haha. This is easy then. slope of the perpendicular line is -2/3 (as m1*m2=-1 in case of perpendicular lines) and it passes through (-6,3). now write the general equation y=mx+c, you know m,y,x. substitute and find c, and put it back into the equation and you have your line.

OpenStudy (anonymous):

huh???

OpenStudy (anonymous):

Didnt get it?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

right now i have a 63

OpenStudy (anonymous):

http://answers.yahoo.com/question/index?qid=20100820235355AAuiJpb This might help you. I think you need to revise your concepts on slopes and straight lines first. I would have explained, but i'm in a hurry atm.